OpenStudy (anonymous):

If (-3, y) lies on the graph of y = 3^x, then y = 1/27 -1 -27

OpenStudy (adi3):

wht do u think

OpenStudy (anonymous):

1/27 ?

OpenStudy (adi3):

How did u do it

OpenStudy (anonymous):

I didn't..lol educated guess

OpenStudy (adi3):

where the graph

OpenStudy (adi3):

where is the graph

OpenStudy (adi3):

do u hve the graph

OpenStudy (anonymous):

There is no graph.

OpenStudy (sepeario):

i'm guessing you have to substitute values for x and y?

OpenStudy (anonymous):

Idk. That's literally all the question says.

OpenStudy (skullpatrol):

y = 3^(-3)

OpenStudy (sepeario):

\[y=3^{-3}\]

OpenStudy (sepeario):

hence y = 1/27

OpenStudy (anonymous):

That's not one of the answer choices.

OpenStudy (sepeario):

haha, but you have to evaluate 3 to the power of -3!

OpenStudy (anonymous):

So it's 1/27 ..? Lol that's what my guess was.

OpenStudy (skullpatrol):

a^(-n) = 1/a^n

OpenStudy (anonymous):

@Sepeario

OpenStudy (sepeario):

yes?

OpenStudy (anonymous):

It is 1/27 ..?

OpenStudy (sepeario):

yes, as @skullpatrol stated due to the evaluation of exponents.

OpenStudy (anonymous):

Thanks y'all! :-)

OpenStudy (skullpatrol):

$$\Huge a^{-n}= \dfrac{1}{a^n}$$

OpenStudy (skullpatrol):

Thanks for asking :)

OpenStudy (phi):

in case it's not obvious: you can invert a number to remove the negative exponent \[ 3^{-3} = \frac{1}{3^3} \] and remember that \(3^3 \) means \(3 \cdot 3 \cdot 3\)

OpenStudy (skullpatrol):

With the "obvious" exception of zero.